OpenStudy (anonymous):

a stone is thrown vertically upward with a speed of 12.5m/s fro the edge of a cliff 75.0m high. How much later does it hit the bottom of the cliff, what is the speed just before hitting? what total distance did it travel?

OpenStudy (anonymous):

i know that: Vi=12.5m/s, Xi=75m, X=-75m, a=g=9.8m/s^2 but i don't know which equations to use to find time,final velocity and total displacement :/

OpenStudy (anonymous):

Xf=0m a=-g=-9.8m/s^2

OpenStudy (anonymous):

okay, so a=-g for the 1st half of the motion when it is thrown upward then afterwards when its going down its a=g because it going down since the ball was thrown at the edge of a cliff..

OpenStudy (anonymous):

Now use \[x _{f}-x_{i}=v_{i}t+at^2/2\]

OpenStudy (anonymous):

No, it is always -g

OpenStudy (anonymous):

how come?

OpenStudy (anonymous):

You see, you are specifying a coordinate system in which the bottom of the cliff is 0, upward direction is +X & the downward direction is -X g=down=negative

OpenStudy (anonymous):

the direction of acceleration never changes. It is only the direction of the velocity that changes. v=>up=+ve =>down=-ve

OpenStudy (anonymous):

oh okay (: thx

OpenStudy (anonymous):

ah! i got it! t=5.39s^^

OpenStudy (anonymous):

do i use v=vi+at to find the velocity before it hit the bottom?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

but i got t=2.2s

OpenStudy (anonymous):

oh wait..i know what i did wrong now my xi and xf where mixed =.=

OpenStudy (anonymous):

no , t=3.48s. Sorry about that.

OpenStudy (anonymous):

its k so v=12.5+(-9.8 x3.48)

OpenStudy (anonymous):

yep.

OpenStudy (anonymous):

and total displacement will just be rearranging v^2=vi^2+2as to solve for s?

OpenStudy (anonymous):

you got it!

OpenStudy (anonymous):

^^ thanks to you!

OpenStudy (anonymous):

Ahh.. but wait. I was blind

OpenStudy (anonymous):

lol okay..

OpenStudy (anonymous):

total distance traveled is what is asked.

OpenStudy (anonymous):

oh okay so how do i find it then? :s

OpenStudy (anonymous):

for that, you have to calculate how high (from the edge of the cliff) the stone traveled.

OpenStudy (anonymous):

okay then add it to the height of the cliff

OpenStudy (anonymous):

if that is 's', total distance traveled=2s+75

OpenStudy (anonymous):

oh okay :) thank you!

OpenStudy (anonymous):

yw:)

OpenStudy (anonymous):

2s=v^2-vi^2/a right?

OpenStudy (anonymous):

No, it is